OpenStudy (drummergirl7):

Starting from the same place, Ryu walks due west and Samantha walks due east. On the number line, 0 represents their starting point, negative values represent the distance traveled west and positive values represent the distance traveled east. All values are given in feet. A number line with a point for Samantha at coordinate 40 and a point for Ryu at coordinate -160. How many feet apart are Ryu and Samantha on the number line? Enter your answer in the box. feet

OpenStudy (drummergirl7):

@StudyGurl14 @awesomedude2078 @AloneS

OpenStudy (drummergirl7):

someone plz help me plzz

OpenStudy (drummergirl7):

i say 9

OpenStudy (studygurl14):

What is 40 + 160?

OpenStudy (studygurl14):

That will be your answer

OpenStudy (drummergirl7):

200

OpenStudy (mathstudent55):

How far is -160 from zero? How far is 40 from zero? Add the amounts.

OpenStudy (mathstudent55):

Also, you should learn this: The distance between any two points on the number line is the absolute value of the difference of their coordinates.

OpenStudy (mathstudent55):

One point has coordinate -160. The other point has coordinate 40. Find the difference between the points and take the absolute value. It does not make a difference which point you use first and which point you use second. A distance is always non-negative, and the absolute value takes care of that. \(|-160 - 40| = |-200 | = 200\) \(|40 - (-160)| = |40 + 160| = |200| = 200\) As you can see, whether you choose 40 or -160 to be first int he subtraction, the answer is still 200.